Reece:

Trevon and Shanice each improved their yards by planting hostas and ivy. (kinds of plants) They bought their supplies from the same store. Trevon spent $130 on 8 hostas and 11 pots of ivy. Shanice spent $70 on 5 hostas and 5 pots of ivy. Find the cost of one hosta and the cost of one pot of ivy.

Reece:

i can answer your qustion

DumbFounded:

Reece this is your question.

Reece:

dumbfounder i can answer your qustion but can you help me

DumbFounded:

oh ok

DumbFounded:

im not good at math problems like these

Reece:

william brocius beter known as curly bill brocius was gunman rustler and outlaw cowboy in the cochisecountry area of arizona

Reece:

thats your answer buddy.

Reece:

@jhonyy9 can u help mee

jhonyy9:

h hostas p pots of ivy 8h +11p = 130 5h +5p = 70 solve for h and p this system

Reece:

ok plz wait

Reece:

let me try and plz don'tleave

Reece:

is it h= 16.25 p=11.81

Reece:

i divided 130 by 11 and 8

jhonyy9:

then you substitute these values inside these equations you get the left side equal right side 5h +5p = 70 5*16,25 +5*11,81 = 70 so this is true ?

jhonyy9:

i think no bc. just 5*16 = 80

Reece:

ok so u wanted me to equal both of the eqautions

jhonyy9:

no 5h +5p = 70 5(h+p)=70 divide both sides by 5 what will get ?

Reece:

so 14

jhonyy9:

h+p = 14 right ?

Reece:

yes thats what i got

jhonyy9:

ok and now there is h+p = 14 => h = 14 -p this h= 14 -p substitute inside the first equation in place of h and solve it for p

Reece:

so if i divide 130 by 5 = 26 right???

jhonyy9:

no you dont need divide 130 by 5

Reece:

so 8(14)+11(p)=

Reece:

so p= .61

jhonyy9:

8h +11p = 130 5h +5p = 70 no 8(14 -p) +11p =130 hence

jhonyy9:

make the calcule and solve it for p

Reece:

so p=6

jhonyy9:

sure ? how you get it ?.

jhonyy9:

64 +66 = 130 correct

Reece:

???

jhonyy9:

ok p = 6 and now calcule the h = ?

jhonyy9:

h = 14 -p so h = ?

Reece:

h= 8 ???

jhonyy9:

h = 14 -6 = 8 h=8 p=6 do you understand it now ?

Reece:

so the hosta is 8 and the pot of ivy is 6 right????

jhonyy9:

exactly perfect

Reece:

ok thank sooo much for being helpful and patient

jhonyy9:

np yw anytime hnd cya

Reece:

bye